OpenStudy (anonymous):

i need to use the equation y = a(x - h)2 + k, my vertex of the parabola reaches 5, -5. I was wondering how would i write the quadratic equation for this ?

OpenStudy (anonymous):

please help someone?

OpenStudy (anonymous):

Not information. The most you could do without another point on the parabola, what you have is y = a(x - 5)^2 - 5. See how there's still and a? You need more info.

OpenStudy (anonymous):

well im not sure what im doing, here ill show u my graph: one sec

OpenStudy (anonymous):

Using the direction, size and vertex of your parabola, think of a quadratic equation in general form, y = a(x - h)2 + k, that would closely match your parabola. Once the graphed parabola matches as close as possible to the image, make note of the quadratic equation, both in general and standard form (y = ax2 + bx + c). -this is what it is asking please help